An amount of $2450 is borrowed for two years at a rate of 6.1% compounded continuously. Payments of $1000 each are made in six months and one year, determine the amount needed to settle the loan at the end of the term.

Please please please show work. I am struggling to figure out why my multiplication with e^x is never coming out calculated correctly.
Answer #1
EAR =[ e^(Annual percentage rate) -1]*100
6.1=(e^(APR%/100)-1)*100
APR% = 5.9212
Future value = present value*(1+ rate)^time
2450 = Present value*(1+0.059212)^2
Present value = 2183.7368
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
5.9212 = ((1+Stated rate%/(2*100))^2-1)*100
Stated rate% = 5.8361
PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 1000*((1-(1+ 5.8361/200)^(-1*2))/(5.8361/200))
PV = 1915.7445

Remaining PV = 2183.7368-1915.7445

=267.9923

Future value = present value*(1+ rate)^time
Future value = 267.9923*(1+0.059212)^2
Future value = 300.67